# Hinge non-Hermitian skin effect in the single-particle properties of a strongly correlated f-electron system

### v1

@@ I Introduction
Quantum systems are often inherently open and subject to dissipative processes. Non-Hermitian operators naturally describe such open quantum systems, where energy, i.e., the eigenvalues of the Hamiltonian, are not necessarily real-valued. These complex-valued band structures offer novel topological features that do not have Hermitian equivalents.Gong et al. (2018); El-Ganainy et al. (2018); Kawabata et al. (2019); Ashida et al. (2020); Bergholtz et al. (2021, 2021); Yokomizo and Murakami (2019); Esaki et al. (2011); Hu and Hughes (2011); Martinez Alvarez et al. (2018) These features include the exceptional point (EP)Heiss (2012); Müller and Rotter (2008); Shen et al. (2018); Miri and Alù (2019); Wang et al. (2022); Budich et al. (2019); Okugawa and Yokoyama (2019); Yoshida et al. (2019); Denner et al. (2021); Wojcik et al. (2020); Yang et al. (2021) and the non-Hermitian skin effect (NHSE).Hatano and Nelson (1996); Yao and Wang (2018); Ezawa (2019); Lee and Thomale (2019); Luo and Zhang (2019); Borgnia et al. (2020); Kawabata et al. (2020a, b); Longhi (2020); Zhang et al. (2020, 2022); Okuma and Sato (2023); Nakamura et al. (2023); Xiujuan Zhang and Chen (2022); Sone et al. (2020); Lin et al. (2023); Kunst et al. (2018); Okuma et al. (2020); Lee (2016); Lieu (2018) The EP represents a band touching with a degeneracy of the eigenstates, which makes the Hamiltonian non-diagonalizable. On the other hand, the NHSE’s extreme sensitivity of eigenvalues and eigenstates originate from non-trivial point-gap topology in the bulk. The localized boundary modes under open boundary conditions are known as skin modes. Non-Hermiticity is not limited to open quantum systems.Wiersig (2022); Minganti et al. (2019); Song et al. (2019); Haga et al. (2021); Ding et al. (2015) It also appears in various other systems, such as photonic crystals,Ozawa et al. (2019); Rüter et al. (2010); Regensburger et al. (2012); Zhen et al. (2015); Zhou et al. (2018); Miri and Alù (2019); Lin et al. (2016); Parto et al. (2021); Özdemir et al. (2019); Zhong et al. (2021); Zhu et al. (2020); Zhu and Gong (2023) electric circuits,Liu et al. (2023, 2021); Helbig et al. (2020); Hofmann et al. (2020); Yoshida et al. (2020a) and correlated systems.Okuma and Sato (2021); Yoshida (2021); Yoshida et al. (2018, 2019, 2020b); Michishita et al. (2020); Michishita and Peters (2020); Kaneshiro et al. (2023); Nagai et al. (2020); Rausch et al. (2021); Crippa et al. (2021); Kaneshiro et al. (2023)
@@
The remainder of this paper is organized as follows. In Sec. II, we introduce the model and methods used. In Sec. III.1, we prove the existence of EPs on the surface of this system. This is followed by the demonstration of the NHSE on the surface for a range of momenta in Sec. III.2. Finally, in Sec. IV, we summarize and conclude this paper.
@@ II.2 Method
where $\omega$ is the frequency, $H_{0}$ is the noninteacting Hamiltonian, $\eta$ is an infinitesimal small, positive adiabatic factor, and $\Sigma(\omega)$ is the self-energy. We note that the self-energy depends on the layer, orbital ($f$ or $c$), and the spin direction. Thus, $H_{\text{eff}}$ is a matrix. Furthermore, this matrix is a non-Hermitian matrix due to the imaginary part of the self-energy. Because the self-energy is frequency-dependent, the effective Hamiltonian also depends on the frequency. In this paper, we focus on the Fermi energy, $\omega=0$, and use $\Sigma(\omega=0)$. The single-particle properties of this system are completely determined by the eigenvalues and eigenstates of this effective Hamiltonian. The single-particle properties at the Fermi energy are thereby mostly given by the eigenvalues with $\text{Re}(E_{i})\approx 0$ because then $\text{Re}(\omega-E_{i})\approx 0$. Thus, we will focus on the eigenvalues of the effective Hamiltonian with a small real part.
@@ III.2 Hinge skin effect induced by surface EPs
For $k_{x}=-0.5$ [see Fig. 5(a)], we see that there are PBC bands connecting the bulk bands for $\text{Re}(E)/|t_{z}|<-0.4$ with the bulk bands $\text{Re}(E)/|t_{z}|>0.4$. These bands clearly form a loop structure, i.e., a point gap opens. We have confirmed that the eigenstates forming this loop structure for $|\text{Re}(E)|/|t_{z}|<0.4$ are located on the $z$ surface and that the bottom and top layers contribute to these eigenstates. The layer dependence of the eigenstates in this region is thereby similar to the eigenstates shown in Fig. 4(b). To confirm the topological point gap, we calculate the winding in Fig. 5(b) for a reference energy ($E_{0}/|t_{z}|=-1.6i$) located inside this point gap. We see that the winding number becomes $-2$ for this reference energy. This winding number indicates a winding of eigenvalues on the top and bottom $z$ surfaces. A winding number of $-2$ is consistent with the fact that these eigenstates are located on the bottom and top surfaces and that both surfaces contribute. Having confirmed the existence of a nontrivial point-gap topology in the energy spectrum, we expect the emergence of the NHSE.Okuma et al. (2020) Analyzing the OBC spectrum for $k_{x}=-0.5$, we see that this loop structure collapses to a single line in the energy spectrum. This sensitivity of the energy spectrum on the boundary conditions is one of the characteristics of the NHSE.
@@ IV Conclusions
We have demonstrated the appearance of the hinge NHSE in the effective Hamiltonian, which describes the single-particle Green’s function. The used model Hamiltonian is derived from a three-dimensional strong topological Kondo insulator under magnetic field and strains. By showing the energy spectrum along a path passing in between one pair of EPs, we have demonstrated the existence of a nontrivial point-gap topology necessary for the emergence of the NHSE. We furthermore have shown that a large number of right eigenstates localize at the right boundary for OBC in the $y$ direction, while the corresponding left eigenstates localize at the left boundary. All these skin modes are localized on the boundaries of the top and bottom $z$ surfaces. Thus, we have demonstrated the emergence of the hinge NHSE in the effective Hamiltonian, describing the single-particle properties of this strongly correlated system.

### v2

@@ I Introduction
Quantum systems are often inherently open and subject to dissipative processes. Non-Hermitian operators naturally describe such open quantum systems, where energy, i.e., the eigenvalues of the Hamiltonian, are not necessarily real-valued. These complex-valued band structures offer novel topological features that do not have Hermitian equivalents.Gong et al. (2018); El-Ganainy et al. (2018); Kawabata et al. (2019); Ashida et al. (2020); Bergholtz et al. (2021, 2021); Yokomizo and Murakami (2019); Esaki et al. (2011); Hu and Hughes (2011); Martinez Alvarez et al. (2018) These features include the exceptional point (EP)Heiss (2012); Müller and Rotter (2008); Shen et al. (2018); Miri and Alù (2019); Wang et al. (2022); Budich et al. (2019); Okugawa and Yokoyama (2019); Yoshida et al. (2019); Denner et al. (2021); Wojcik et al. (2020); Yang et al. (2021) and the non-Hermitian skin effect (NHSE).Hatano and Nelson (1996); Yao and Wang (2018); Ezawa (2019); Lee and Thomale (2019); Luo and Zhang (2019); Borgnia et al. (2020); Kawabata et al. (2020a, b); Longhi (2020); Zhang et al. (2020, 2022); Okuma and Sato (2023); Nakamura et al. (2023); Xiujuan Zhang and Chen (2022); Sone et al. (2020); Lin et al. (2023); Kunst et al. (2018); Okuma et al. (2020); Lee (2016); Lieu (2018) The EP represents a band touching with a degeneracy of the eigenstates, which makes the Hamiltonian non-diagonalizable. On the other hand, the NHSE’s extreme sensitivity of eigenvalues and eigenstates originate from non-trivial point-gap topology in the bulk. The localized boundary modes under open boundary conditions are known as skin modes. Non-Hermiticity is not limited to open quantum systems.Wiersig (2022); Minganti et al. (2019); Song et al. (2019); Haga et al. (2021); Ding et al. (2015) It also appears in various other systems, such as photonic crystals,Ozawa et al. (2019); Rüter et al. (2010); Regensburger et al. (2012); Zhen et al. (2015); Zhou et al. (2018); Miri and Alù (2019); Lin et al. (2016); Parto et al. (2021); Özdemir et al. (2019); Zhong et al. (2021); Zhu et al. (2020); Zhu and Gong (2023) electric circuits,Liu et al. (2023, 2021); Helbig et al. (2020); Hofmann et al. (2020); Yoshida et al. (2020a); Halder et al. (2024) and correlated systems.Okuma and Sato (2021); Yoshida (2021); Yoshida et al. (2018, 2019, 2020b); Michishita et al. (2020); Michishita and Peters (2020); Kaneshiro et al. (2023); Nagai et al. (2020); Rausch et al. (2021); Crippa et al. (2021); Kaneshiro et al. (2023)
@@
The remainder of this paper is organized as follows. In Sec. II, we introduce the model and methods used. In Sec. III.1, we prove the existence of EPs on the surface of this system. This is followed by the demonstration of the NHSE on the surface for a range of momenta in Sec. III.2. Finally, in Sec. IV, we summarize and conclude this paper. Appendices are devoted to the analysis of the $T=0$ case and the non-Hermitian Hamiltonian under open boundary conditions in all directions.
@@ II.2 Method
where $\omega$ is the frequency, $H_{0}$ is the noninteacting Hamiltonian, $\eta$ is an infinitesimal small, positive adiabatic factor, and $\Sigma(\omega)$ is the self-energy. We note that the self-energy depends on the layer, orbital ($f$ or $c$), and the spin direction. Thus, $H_{\text{eff}}$ is a matrix. Furthermore, this matrix is a non-Hermitian matrix due to the imaginary part of the self-energy. We stress that correlations play an essential role in the non-Hermiticity of $H_{\text{eff}}$ in this heavy-fermion system. Because the self-energy is frequency-dependent, the effective Hamiltonian also depends on the frequency. In this paper, we focus on the Fermi energy, $\omega=0$, and use $\Sigma(\omega=0)$. The single-particle properties of this system are completely determined by the eigenvalues and eigenstates of this effective Hamiltonian. The single-particle properties at the Fermi energy are thereby mostly given by the eigenvalues with $\text{Re}(E_{i})\approx 0$ because then $\text{Re}(\omega-E_{i})\approx 0$. Thus, we will focus on the eigenvalues of the effective Hamiltonian with a small real part.
@@ III.2 Hinge skin effect induced by surface EPs
For $k_{x}=-0.5$ [see Fig. 5(a)], we see that there are PBC bands connecting the bulk bands for $\text{Re}(E)/|t_{z}|<-0.4$ with the bulk bands $\text{Re}(E)/|t_{z}|>0.4$. These bands clearly form a loop structure, i.e., a point gap opens. We have confirmed that the eigenstates forming this loop structure for $|\text{Re}(E)|/|t_{z}|<0.4$ are located on the $z$ surface and that the bottom and top layers contribute to these eigenstates. The layer dependence of the eigenstates in this region is thereby similar to the eigenstates shown in Fig. 4(b). To confirm the topological point gap, we calculate the winding in Fig. 5(b) for a reference energy ($E_{0}/|t_{z}|=-1.6i$) located inside this point gap. We see that the winding number becomes $-2$ for this reference energy. This winding number indicates a winding of eigenvalues on the top and bottom $z$ surfaces. We note that the effective Hamiltonian fulfills reflection symmetry in the $z$-direction. A winding number of $-2$ is thus consistent with the fact that these eigenstates are located on the bottom and top surfaces and that both surfaces contribute. Having confirmed the existence of a nontrivial point-gap topology in the energy spectrum, we expect the emergence of the NHSE.Okuma et al. (2020) Analyzing the OBC spectrum for $k_{x}=-0.5$, we see that this loop structure collapses to a single line in the energy spectrum. This sensitivity of the energy spectrum on the boundary conditions is one of the characteristics of the NHSE.
@@ IV Conclusions
We have demonstrated the appearance of the hinge NHSE in the effective Hamiltonian, which describes the single-particle Green’s function. The used model Hamiltonian is derived from a three-dimensional strong topological Kondo insulator under magnetic field and strains. By showing the energy spectrum along a path passing in between one pair of EPs, we have demonstrated the existence of a nontrivial point-gap topology necessary for the emergence of the NHSE. We furthermore have shown that a large number of right eigenstates localize at the right boundary for OBC in the $y$ direction and PBC in the $x$ direction, while the corresponding left eigenstates localize at the left boundary. All these skin modes are localized on the boundaries of the top and bottom $z$ surfaces. Thus, we have demonstrated the emergence of the hinge NHSE in the effective Hamiltonian, describing the single-particle properties of this strongly correlated system.
Furthermore, in Appendix B, we observe the existence of localized eigenstates of the effective Hamiltonian under OBC in all directions. These eigenstates also localize along the hinges of the bottom and top surfaces or at the corners of the system. However, the topological origin of these localized modes is left as a future work.
@@ Appendix A Hermitian boundary states
Appendix B Open boundary conditions
Figure 9: Energy spectrum and one right eigenstate in the system with OBC in all directions. We diagonalize the effective Hamiltonian in the rotated geometry corresponding to Fig. 1(c). Panel (a) shows the imaginary part over the real part of the energy of the effective Hamiltonian. In panel (b), the color represents the amplitude of the right eigenstate [indicated by the red color in panel (a)] over the position in the system.
Figure 10: Right ($\Psi_{R}$) and left eigenstates ($\Psi_{L}$) in the system with OBC in all directions. Panels (a) and (b) show the $f$-orbital, spin-up component of the right and left eigenstate [indicated by the red color in Fig. 9(a)]. The color indicates the amplitude over the position of the atom in the lattice. Panels (c) and (d) show the $f$-orbital, spin-down component of the same right and left eigenstate.
We analyze the effective Hamiltonian in a system with OBC in all directions. We diagonalize the effective Hamiltonian (including the self-energy) in the rotated geometry, corresponding to Fig. 1(c). The size of the system in all three directions is chosen as $(L_{x},L_{y},L_{z})=(10,10,5)$. We note that this results in a Hilbert space dimension of $4000$, which can still be completely diagonalized. Here, we make the simplification that the self-energy only depends on the layer index in the $z$-direction, corresponding to a similar situation as in the previous sections. Then, the imaginary parts of the bottom and top layers are again larger than the imaginary parts of the middle layers, which would lead to the emergence of exceptional points on the bottom and top surfaces in a system with PBC.
In the main text, we showed the emergence of the non-Hermitian skin effect in the $y$-direction using PBC in the $x$-direction. From the position of the exceptional points in Fig. 1(c), it is clear that there is also a skin effect in the $x$-direction when using PBC in the $y$-direction. Thus, the emergence of a non-Hermitian skin effect on the bottom and top layers in the $x$- and $y$-direction seems possible. To confirm this, we show the energy spectrum (imaginary part over real part) of the fully open system in Fig. 9(a). Comparing these energies with Fig. 5, we see that there are eigenstates of the fully open system in the same energy region where we previously observed the non-Hermitian skin effect. In Fig. 9(b), we show the amplitude of a right eigenstate [indicated by the red color in Fig. 9(a)] as a color over the position of the atom in the lattice. We see that this eigenstate is localized at the corners of the bottom and top surfaces. This is a corner state localizing simultaneously in the $x$- and $y$-directions of the surface layers. Contrary to the results in the main text, where the skin effect led to a localization at only one boundary, we now see a localization at two opposite boundaries. We note that many eigenstates with Re(E)$\approx 0$ show a similar behavior. The localization at two opposite boundaries may be understood by the fact that the original multiband Hamiltonian fulfills time-reversal symmetry. The non-Hermitian skin effect in systems with time-reversal symmetry leads to degenerate skin modes localized at opposite boundaries. Although, in the current system, the time-reversal symmetry is broken by a small magnetic field, the breaking of the symmetry is not very strong. Thus, in the fully open system, we can expect a mixing of eigenstates localized at opposite boundaries. To justify this hypothesis, we show the left and right eigenstates of the $f$-orbital with spin-up in Fig. 10(a,b) and the left and right eigenstates of the $f$-orbital with spin-down in Fig. 10(c,d). Although time-reversal symmetry is broken in this effective Hamiltonian, we can nevertheless expect that the spin-up and spin-down components behave approximately in the opposite way. We immediately see that, as expected for skin modes, these eigenstates localize only in one corner of the bottom and top surfaces. Furthermore, the localization of the right and left eigenstates is opposite, which confirms this fundamental property of skin modes. Finally, we see that the localization of the spin-up component and the spin-down component are also opposite. As the full eigenstate shown in Fig. 9(b) consists of spin-up and spin-down components, this state is localized at two opposite corners of the bottom and surface layers. Based on this argument, we believe that the skin modes of this multiband system are localized in two opposite directions on the bottom and top layers.
Thus, the obtained numerical data indicates the existence of skin modes that are localized at the hinges or corners of the bottom and top surfaces under OBC in all directions. However, we stress that the topological characterization of the skin modes performed in the main text is only valid for a system with mixed boundary conditions (OBC in the $x$- and $z$-directions and PBC in the $y$-direction). The topological origin of the skin modes in the fully open system is left as a future work.